Project

General

Profile

Actions

Physics Shift Report

Start Date: 09-29-2017
End Date: 10-03-2017
Start Run: 9773
End Run: 9841
Physics shifter: Tanaz A. Mohayai

Run Number Purpose
9773 - 9779 Scattering study, LH2 full - Cooling Channel magnets OFF
9780 pulser run to test flip mode settings in the cooling channel magnets
9781 - 9804 Emittance reduction - LH2 full - Flip mode
9805 - 9821 Field-on scattering and magnet alignment
9821 - 9834 Emittance reduction - LH2 full - Flip mode (with new cooling channel settings)

Elog, CDB, and Reconstruction

Elog entries are ok.

CDB uploaded ok.

Online reconstruction is ok.

Note the TOF1 reconstruction from run 9834:

Run 9834 input beam setting is 10-140 (task H28c in https://micewww.pp.rl.ac.uk/projects/operations/wiki/RunPlan20171002)
This was not observed in the other 3-140 or 6-140 settings. D. Rajaram is re-starting on-rec to see if it fixes it. The pattern we observe in the above TOF1 spacepoint is not observed in TOF1 raw PMT plot (TofReduce_tof_pm0pln1.png).
Update: re-starting on-rec fixed the problem:

Updated by Mohayai, Tanaz Angelina almost 5 years ago · 12 revisions